OpenStudy (anonymous):

Cylinder A has a radius of 1 m and a height of 4 m. Cylinder B has a radius of 2 m and a height of 4 m. Find the ratio of the volume of cylinder A to the volume of cylinder B.

OpenStudy (anonymous):

s1=pi*r^2*h=pi*1*4=4pi s2=pi*r^2*h=pi*2^2*4=16pi s1/s2=4pi/16pi=1/4

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

welcome